# Generalizable Trajectory Prediction via Inverse Reinforcement Learning with Mamba-Graph Architecture

### v1

@@ III Evaluation and Results
Datesets. Our experimental evaluation leverages two large-scale trajectory datasets—the rounD [20] and inD [21] datasets—which comprise vehicular movement patterns captured at 25Hz across diverse German roundabout and intersection scenarios(Fig. 3). These datasets, having undergone preprocessing procedures outlined in [10], were subsequently downsampled by a factor of 5 to achieve a temporal resolution of 0.2s. The observation window was deliberately constrained to 3s ($L$ = 15 timesteps), while the prediction horizon was extended to 5s ($N$ = 40 timesteps). Through this meticulous curation process, the final datasets yield 29,248 and 3,738 high-fidelity samples for rounD and inD, each sample representing a complete spatiotemporal graph sequence. To ensure robust model development and evaluation, the data was partitioned in a stratified manner, with 80% allocated for training, 10% for validation, and the remaining 10% reserved for testing purposes.
Figure 3: Environment and trajectory notation of urban intersection scenario at Heckstrasse.
Metrics The algorithm’s predictive performance was evaluated using - distinct metrics (Tab. I), each carefully selected to provide comprehensive assessment criteria. Note that the definition of ADE, FDE, MR, APDE and CR follows [10].
TABLE I: Metrics Details
Metrics Description
ADE
Average Displacement Error: The average Euclidean
distance in meters between the predicted and
ground truth trajectories.
FDE
Final Displacement Error: The Euclidean distance
in meters between the predicted and ground truth
trajectories at the final time step.
APDE
Average Path Displacement Error: The average minimum
L2-norm between thepredicted positions and
ground truth.
MR
Miss Rate: The percentage of predictions where
the predicted trajectory at the final time step
is not within 2.0 m from the ground truth.
CR
Crash Rate: the proportion of predicted trajectories
that result in collisions with obstacles or other
agents under simulated scenarios.
TABLE II: ROUND DATASET VEHICULAR TRAJECTORY PREDICTION METRICS
metric ADE FDE MR APDE
@@
TABLE III: VEHICULAR TRAJECTORY PREDICTION METRICS ON CROSS-SCENARIO ADAPTABILITY
metric ADE FDE MR APDE CR
inD(seen)
@@
TABLE IV: ABLATION EXPERIMENT RESULTS
@@
$\textbf{rounD}^{*}$ $\text{H}_{1}$ 4.72 10.27 0.94 2.42 0.00069
@@ III-A Results Analysis on rounD
The quantitative results are presented in Tab. II. In contrast to Mixed Mamba [11]—a contemporaneous approach that similarly incorporates the Mamba module—the cited method purportedly achieves superior accuracy on the highD dataset compared to the rounD dataset. Our method achieves state-of-the-art performance across key metrics—including average displacement error , final displacement error and average path displacement error, e.g., FDE of 2.78, representing an 8% reduction compared to the Mixed Mamba, thereby demonstrating its robustness in handling high-variability driving behaviors. The discrepancy of Mixed Mamba performance can be attributed to the inherent simplicity of highway trajectory prediction, where vehicular motion patterns are predominantly linear. Furthermore, the highD dataset’s extensive scale likely augments the Mamba module’s performance. However, the aforementioned work fails to extend Mamba’s advantages to roundabout trajectory prediction, a domain that presents significantly greater challenges due to the non-linear and dynamic nature of vehicular interactions. The result shows our proposed method addresses these limitations with a maximum cumulative reward objective. This integration enables our framework to excel in complex urban roundabout scenarios, such as these in Aachen and Alsdorf.
@@ III-B The Effectiveness of Maxent IRL
To rigorously evaluate the generalization capability and Cross-Scenario Adaptability (CSA) of our proposed methodology, we conducted comprehensive experiments involving training on the inD dataset while validating performance across both inD and rounD datasets, with quantitative results tabulated in Tab. III. This experimental paradigm was deliberately designed to probe the model’s transferability between ostensibly similar yet distinct driving environments - while both datasets exhibit comparable microscopic behaviors such as car-following, lane-changing, and turning maneuvers within confined spatial contexts, they diverge in road geometries and traffic dynamics between intersections and roundabouts.
@@
with $M_{i}$ being the $i$-th evaluation metric (e.g., ADE, FDE). The terms with “$\max$” and “$\min$” subscripts denote the maximum and minimum values of certain scenario respectively across all evaluated methods, respectively, while $\alpha,\beta$ represents the weighting coefficient. Intuitively, a higher CSA value indicates superior cross-scenario adaptability.
@@
The qualitative results are shown in Fig. 4. For interpretability, we set $\alpha=0.8$, $\beta=0.05$ to account for the substantially larger magnitude of performance degradation compared to inter-method variance on individual metrics. Notably, although the CSA values for MR are comparatively lower across all methods—reflecting the inherent challenge of behavioral prediction in unseen scenarios—our approach still demonstrates state-of-the-art performance within this category. More significantly, our method outperforms all baselines in the remaining metrics, achieving approximately 2 times the CSA scores of the second-best method in both APDE and FDE.
@@
By virtue of employing Maxent IRL, which endows the learned policies with substantially enhanced diversity, our method demonstrates superior performance in the inD scenario (Tab. III). Specifically, the proposed approach achieves significantly lower average displacement error and final displacement error compared to the apprenticeship learning method (AppL [16]) that adheres to the maximum margin theory framework. This empirical result substantiates our framework’s enhanced capacity for holistically approximating the underlying distribution of human driving trajectories, demonstrating more faithful reproduction of naturalistic behavioral patterns compared to AppL approaches.
@@
Ablation Study. The ablation experiments for the proposed method were conducted on the inD and rounD datasets, with the results summarized in Tab. IV. Notably, $\textit{rounD}^{*}$ denotes the cross-scenario evaluation where the model is trained exclusively on urban intersection scenarios and is tested on roundabout environments. The extend of degradation of mmTransformer and GNN-RNN methods in transfer scenarios may be attributed to model complexity and GNN-based environmental encoding mechanisms. Generally speaking, more sophisticated temporal modeling architectures, while exhibiting stronger modeling capacity, tend to demonstrate correspondingly poorer generalization capabilities in novel scenarios. Further analysis on the tabulated data reveals that both the MaxEnt IRL module and the GATs module contribute to trajectory prediction accuracy, with the latter exhibiting even more pronounced enhancement. This phenomenon can be attributed to the fact that the graph-based architecture, introduces additional heterogeneous graph data encompassing information that the IRL method doesn’t contain. However, the IRL component does play a pivotal role in bolstering the model’s generalizability across scenarios: in the $\textit{rounD}^{*}$ column in Tab. IV, the absence of the IRL module precipitates an 83% deterioration in ADE and a more than twofold decline in FDE.
@@ IV Conclusion
We proposed the Environment-Aware Mamba Trajectory Predictor with MaxEnt IRL, a novel framework that integrates Mamba-based sequence modeling with IRL to capture human driver behavior. We developed an encoder-decoder architecture combining Mamba blocks for efficient long-sequence dependency modeling and GATs for encoding spatial interactions among traffic agents. The MaxEnt IRL module further enhanced the framework’s adaptability by inferring diverse driver reward functions from heterogeneous demonstrations, enabling human-like and scenario-aware trajectory generation. Experiments on urban intersections and roundabouts demonstrated that our method achieves state-of-the-art prediction accuracy while exhibiting strong generalization to unseen scenarios, owing to its ability to learn transferable reward structures from human data.
Future research directions could explore explicitly deriving agent policies through the learned reward function, which not only has the potential to further enhance the method’s cross-scenario adaptability but also enables more stable training through techniques like importance sampling for approximating the partition function $Z$ with $\pi^{*}$, broadening the practical applicability in real-world autonomous driving systems.

### v2

@@ III Evaluation and Results
Datesets. Our experimental evaluation leverages four large-scale trajectory datasets—the rounD [12], inD [3], highD [11] and exiD [16] datasets—which comprise vehicular movement patterns captured at 25Hz across diverse German roundabout, intersectio (Fig. 3), highway entries and exits scenarios. These datasets, having undergone preprocessing procedures outlined in [22], are subsequently downsampled by a factor of 5 to achieve a temporal resolution of 0.2s. The observation window and prediction horizon are set to 3s (LL=15) and 5s (NN=40) for the inD and rounD datasets, and to 2s and 5s for the highD and exiD datasets, respectively. The data is split into a stratified 80-10-10 partition for training, validation, and testing.
Although IRL is potentially sensitive to the quality of demonstrations, we made no attempts to filter or curate the data to avoid inconsistent or unsafe behaviors. Note that in all cross-scenario testing described herein, models are trained only on source domain data and evaluated solely on target domain data. Highway scenarios (including entries and exits) serve as OOD test environments, characterized by fundamentally distinct probability distributions of lane-changing, turning, acceleration, and braking behaviors compared to the source domain of urban intersections and roundabouts, as substantiated by the baseline evaluation results in OOD test (Tab. V).
Figure 3: Environment and trajectory notation of urban intersection scenario at Heckstrasse [3].
Metrics The algorithm’s predictive performance was evaluated using five key metrics: Average Displacement Error (ADE), Final Displacement Error (FDE), Average Path Displacement Error (APDE), Miss Rate (MR), and Crash Rate (CR). Their definitions follow [22].
TABLE I: RounD Dataset Vehicular Trajectory Prediction Metrics
Model ADE FDE MR APDE
@@
TABLE II: Vehicular Trajectory Prediction Metrics on Cross-Scenario Adaptability
Model ADE FDE MR APDE CR
inD (Seen Scenario)
@@
TABLE III: Ablation Experiment Results
@@
$\textit{rounD}^{*}$ $\text{H}_{1}$ 4.72 10.27 0.94 2.42 0.00069
@@ III-A Results Analysis on rounD
The quantitative results are presented in Tab. I. In contrast to Mixed Mamba [23]—a contemporaneous approach that similarly incorporates the Mamba module—the cited method purportedly achieves superior accuracy on the highD dataset compared to the rounD dataset. Our method achieves state-of-the-art performance across key metrics—including ADE, FDE and APDE, e.g., FDE of 2.78, representing an 8% reduction compared to the Mixed Mamba, thereby demonstrating its robustness in handling high-variability driving behaviors. The discrepancy of Mixed Mamba performance can be attributed to the simplicity of highway trajectory prediction, where vehicular motion patterns are predominantly linear. Furthermore, the highD dataset’s extensive scale likely augments the Mamba module’s performance. However, the aforementioned work fails to extend Mamba’s advantages to roundabout trajectory prediction, a domain that presents greater challenges due to the non-linear and dynamic nature of vehicular interactions. The result shows our proposed method addresses these limitations with a maximum cumulative reward objective. This integration enables our framework to excel in complex urban roundabout scenarios, such as these in Aachen and Alsdorf.
@@ III-B The Effectiveness of MaxEnt IRL
To evaluate the generalization capability and Cross-Scenario Adaptability (CSA) of our proposed methodology, we conducted experiments involving training on the inD dataset while validating performance across both inD and rounD datasets, with quantitative results tabulated in Tab. II. This experimental paradigm was deliberately designed to probe the model’s transferability between ostensibly similar yet distinct driving environments - while both datasets exhibit comparable microscopic behaviors such as car-following, lane-changing, and turning maneuvers within confined spatial contexts, they diverge in road geometries and traffic dynamics between intersections and roundabouts.
@@
with $M_{i}$ being the $i$-th evaluation metric (e.g., ADE, FDE). The terms with $\max$” and $\min$” subscripts denote the maximum and minimum values of certain scenario respectively across all evaluated methods, respectively, while $\alpha,\beta$ represents the weighting coefficient. Intuitively, a higher CSA value indicates superior cross-scenario adaptability.
@@
The qualitative results are shown in Fig. 4. Notably, although the CSA values for MR are comparatively lower across all methods—reflecting the inherent challenge of behavioral prediction in unseen scenarios—our approach still demonstrates state-of-the-art performance within this category. More significantly, our method outperforms all baselines in the remaining metrics, achieving 2.3 times the CSA scores (in terms of APDE) of the second-best method.
@@
By virtue of employing MaxEnt IRL, which endows the learned policies with enhanced diversity, our method demonstrates superior performance in the inD scenario (Tab. II). Specifically, the proposed approach achieves significantly lower ADE and FDE compared to the apprenticeship learning method (AppL [4]) that adheres to the maximum margin theory framework. This empirical result substantiates our framework’s enhanced capacity for holistically approximating the underlying distribution of human driving trajectories, demonstrating more faithful reproduction of naturalistic behavioral patterns compared to AppL approaches.
TABLE IV: Computational Efficiency Comparison of Different Models
Time/Param Memory/Param
Model (s/1000 params) (GB/1000 params)
Mamba 0.0184 0.0130
Transformer 0.2255 0.1452
GNN 0.0737 0.2664
Note: The Mamba model demonstrates the highest computational efficiency among the three models in terms of training memory usage relative to model size and inference time relative to model size. To avoid differences in inference time caused by varying pipelines, we only record the pure model inference time. Batch size is set to 128.
@@
Ablation Study. The ablation experiments for the proposed method were conducted on the inD and rounD datasets, with the results summarized in Tab. III. Notably, $\textit{rounD}^{*}$ denotes the cross-scenario evaluation. The Mamba module serves as the backbone component of the predictor and was not included in the ablation study. The claim of Mamba’s exceptional inference efficiency is supported by the results in Tab. IV. Further analysis on the tabulated data reveals that both the MaxEnt IRL module and the GATs module contribute to trajectory prediction accuracy, with the latter exhibiting even more pronounced enhancement. This phenomenon can be attributed to the fact that the graph-based architecture, introduces additional heterogeneous graph data encompassing information that the IRL method doesn’t contain. However, the IRL component does play a pivotal role in bolstering the model’s generalizability: in the $\textit{rounD}^{*}$ column in Tab. III, the absence of the IRL module precipitates an 83% deterioration in ADE and a more than twofold decline in FDE.
III-C Out-of-Distribution Scenarios Test
TABLE V: OOD Scenarios Test Results
Dataset Method ADE APDE FDE
highD Baseline 638.01 580.1 678.35
+ Policy 290.81 228.8 377.41
LoRA 233.68 177.64 259.68
exiD Baseline 243.95 198.59 285.2
+ Policy 89.89 43.41 112.89
LoRA 99.04 52.69 118.94
The OOD adaptability experiment on the highD and exiD dataset validates our method’s capability to generalize in unfamiliar OOD scenarios. Based on our graph based Mamba predictor, we conduct two sets of experiments: a baseline model augmented with the explicit policy extension (Fig. 6), and baseline model finetuned accordingly with highD/exiD ground truth data using Low-Rank Adaptation (LoRA) [9]—a parameter-efficient fine-tuning (PEFT) method (Fig. 5). In the highway scenario, where kinematic models differ substantially from those in urban intersections and roundabouts, the policy extension (’+Policy’ in Tab. V) improves the baseline performance but slightly underperforms compared to the LoRA-based fine-tuning. In contrast, on the exiD dataset—also a highway environment but with a higher frequency of merging and exiting behaviors—the policy extension demonstrates marginally better adaptability than LoRA.
Figure 5: Metrics on validation set obtained from fine-tuning on 100 random instances of the highD dataset, exhibiting convergence within 6 epochs. Following experimental results, the input projection, time parameter projection, and output projection modules in the Mamba block are selected as the optimal targets for LoRA fine-tuning with a parameter compression rate of 3.91%.
Figure 6: Validation metrics obtained during policy learning using the TD3 algorithm, showing convergence within 80 epochs. The policy delay for TD3 is set to 3. The policy is trained once in the source domain and can be directly applied to any target domain (e.g., highD or exiD)
The cross-scenario adaptability of our method in analogous and OOD scenario is further accentuated by the trade-off in Fig. 7: increased model complexity (larger parameters) improves in-domain accuracy but often leads to degradation of cross-scenario generalization. This result underscores the superiority of our approach for leveraging the degradation effect of introducing mamba architecture (with more parameters) and still preserves cross-scenario adaptability.
Figure 7: An illustration of the accuracy-parameter-degradation trade-off in cross-scenario adaptation (urban intersection to highway). FDE (vertical axis) is plotted against model parameter size (horizontal axis, in thousands), with the color intensity of each point representing its degree of performance degradation - darker colors represent more severe degradation.
@@ IV Conclusion
We proposed the environment-aware Mamba trajectory predictor with MaxEnt IRL, a novel framework that integrates Mamba-based sequence modeling with IRL to capture human driver behavior. We developed an encoder-decoder architecture combining Mamba blocks for efficient long-sequence dependency modeling and GATs for encoding spatial interactions among traffic agents. The MaxEnt IRL module and policy extension further enhanced the framework’s adaptability by inferring diverse driver reward functions from heterogeneous demonstrations, enabling human-like and scenario-aware trajectory generation. Experiments on urban intersections, roundabouts and highways demonstrated that our method achieves state-of-the-art prediction accuracy while exhibiting strong generalization to unseen scenarios, owing to its ability to learn transferable reward structures from human data.
Future research may explore leveraging agent policies to stabilize training through techniques like importance sampling for approximating the partition function $Z$ with $\pi^{*}$, and better integrate long- and short-horizon predictions to improve generalization, thereby broadening the practical applicability in real-world autonomous driving systems.